What is the amount of energy needed to increase the temperature of 2.5 mole of water from 20oC to 100oC. Cpwater = 9R
Answer #1

We know Heat required to rais the temperature (q) = m*Cp*T

m=moles= 2.5 mole

Cp = 9R = 8.314J/molK * 9 = 74.826 J/molK .

T = 100-20 = 80oC = 80 K

q = 2.5*74.826*80 = 14965.2 J = 14.9652 kJ

thus to increase the temprature of 2.5mol of water we need 14.9652 kJ of energy.